# Adversaries With Incentives: A Strategic Alternative to Adversarial Robustness

### v1

@@ 6.3 Worst-case strategic robustness
If our knowledge of the opponent’s utility is only coarse (and not specific), then a natural approach is to seek robustness to a set of plausible utilities $U$ (rather than focus on a single $u$). Motivated by our previous results, our goal in this section is to explore strategic training when all we assume is that the opponent benefits either from attacks that preserve, or that disrupt, label semantics.
@@
In addition to CIFAR-10, here we also experiment with the GTSRB road signs dataset, in which we partition road sign classes according to their type. We consider three architectures: VGG, ResNet18, and vision transformer (ViT). For adversarial training we implement attacks using PGD [28], and for strategic training we use our approach from Sec. 5. Since we are interested in strategic robustness, we evaluate the performance of each method as its accuracy against the worst-case strategic opponent in the relevant incentive set. i.e., $U^{\mathrm{sem}}$ for semantic and $U^{\mathrm{a{\text{-}}sem}}$ for anti-semantic. Appendix B includes further details on data, optimization, and evaluation.
@@ Results.
For both semantic and anti-semantic settings, Table 1 compares models trained using clean, strategic, and adversarial training in terms of their clean, strategic, and adversarial accuracies. As can be seen, strategic training provides significant gains over adversarial training against worst-case strategic opponents in all cases, with a more pronounced effect in the anti-semantic setting. This is made apparent in Table 7, which shows the relative gain of strategic training, measured as the percentage of attacks it is able to prevent out of the successful attacks against an adversarial model:
%$\%\mathrm{deflected}=(\mathrm{strat}(f_{\mathrm{str}})-\mathrm{strat}(f_{\mathrm{adv}}))/(\mathrm{clean}(f_{\mathrm{cln}})-\mathrm{strat}(f_{\mathrm{adv}}))$ (11)
@@ Appendix A Broader impact
The field of adversarial learning aims to enable safer deployment of learned models and to make the use of machine learning more secure in real-world applications. The concern regarding attacks on models is concrete, and the possible risks such attacks entail span many domains [24]. This underlies the many efforts that are and have been put forth, both in research and in industry, to advance the development of methods for adversarially robust learning.
@@
Our approach to robust learning stems from the observation that in reality, learned models will be susceptible to the influence of the agents who interact with them; and whereas adversarial learning prepares to encounter malicious agents, our main thesis is that even non-benign agents are more likely to promote their own interests than to sabotage learning efforts. As such, we view our work as contributing an economic perspective to how robust learning should be designed, this by acknowledging and modeling the possible utilities of the agents that learning may encounter.
@@
However, as in any domain or task in which agents are modeled, our approach necessitates making assumptions—even if mild—on the nature of their possible incentives. As these are intended to relax the demanding restrictions of adversarial learning, the main limitation of our approach is that the more precisely we model incentives, the more we risk their misspecification. This is an inherent limitation of any method for robust learning: once we commit to an uncertainty set, all guarantees are with respect to this set only. In our work we empirically explore (to some extent) the implication of misspecification (Sec. 6.2). Results are fairly encouraging, but safe deployment of strategically trained models clearly requires much further investigation of possible failure modes.
@@
Adversarial learning faces similar limitations: it commits to a certain threat model (e.g., $L_{\infty}$ with a certain budget), and provides robustness against such attacks—but this protection does not generalize to others (e.g., $L_{2}$, or a larger budget) [19, 2]. Nonetheless, in relation to our approach, adversarial learning is indeed ‘safer’, but this is only because it seeks to defend against all possible opponents (i.e, ‘everything’). Our key point in this paper is precisely that safety that is granted through conservativity comes at a cost—for example, in reduced clean performance—and that if we have established knowledge or strong beliefs regarding the opponent’s possible incentives, then it can be highly beneficial to use them as inductive bias for strategic learning.
@@ B.1 Data and preprocessing
  • CIFAR-10333https://www.cs.toronto.edu/~kriz/cifar.html [23]: This dataset include $60,000$ natural images of size $32\times 32\times 3$. We use the standard split of $50,000$ train and $10,000$ test examples. There are $10$ classes, 4 of which encode vehicles, and 6 of which encode animals. We use this partition as our definition of label semantics.
  • GTSRB [^4] https://benchmark.ini.rub.de [21]: This dataset includes $\sim 50,000$ images of road signs of variable sizes. We use the standard split of $39,209$ train and $12,630$ test images For consistency we resize the images to $32\times 32\times 3$. There are $43$ classes, which we partition according to the shape of the sign—either circular or triangular, which roughly corresponds to road sign types: prohibitory and mandatory signs vs. warning signs.
@@ B.2 Architectures
We experiment with three different model architectures: (i) VGG [39], (ii) ResNet18 [18], and (ii) ViT [11]. The former two are convolution-based, for which we use standard architectures. The latter is transformer-based, and for this we use an implementation adjusted to smaller resolution images. [^5] https://github.com/omihub777/ViT-CIFAR
@@ B.4 Training and hyperparameters
As we discuss in Sec. 5.1, we found that strategic training benefits (mostly in terms of clean accuracy) from adding subtle noise to the choice of utility at each step. Accordingly, at each batch, for each class $y$, with probability $1-\epsilon$ we use the utility $u_{y\cdot}$ defined in the objective, and with probability $\epsilon$ we draw a target label ${\bar{y}}\in[K]$ uniformly at random. Throughout all experiments we use $\epsilon=0.1$.

### v3

@@ 6.3 Worst-case strategic robustness
Motivated by our previous results, in this section we consider strategic learning that seeks robustness against the worst-case opponent within a set of plausible utilities $U$. Here we present results for $U$ that consists of either all semantic or anti-semantic utilities. This means that all we will assume is that the opponent benefits either from attacks that preserve, or that disrupt, label semantics. Our choice is based on the observation that adversarial attacks tend to preserve semantics; hence, semantic and anti-semantic serve us as representative ‘hard’ and ‘easy’ strategic sets, respectively. In Appendix C.4 we present additional experiments for preference-order utilities, for which results exhibit similar trends.
@@
In addition to CIFAR-10, here we also experiment with: (i) the GTSRB road signs dataset (43 classes), in which we partition road sign classes according to their type, and (ii) CIFAR-100 (100 classes), reported in Appendix C.5. We consider three architectures: VGG, ResNet18, and vision transformer (ViT). For adversarial training we implement attacks using PGD (Madry et al., 2018), and for strategic training we use our approach from Sec. 5. Since we are interested in strategic robustness, we evaluate the performance of each method as its accuracy against the worst-case strategic opponent in the relevant incentive set. i.e., $U^{\mathrm{sem}}$ for semantic and $U^{\mathrm{a{\text{-}}sem}}$ for anti-semantic. Appendix B includes further details on data, optimization, and evaluation. Appendix C.3 includes sensitivity tests against a stronger opponent who knows the uncertainty set.
@@ Results.
For both semantic and anti-semantic settings, Table 1 compares models trained using clean, strategic, and adversarial training in terms of their clean, strategic, and adversarial accuracies. As can be seen, strategic training provides significant gains over adversarial training against worst-case strategic opponents in all cases, with a more pronounced effect in the anti-semantic setting. Strategic training also exhibits generally improved clean accuracy. Importantly, strategic training also provides a reasonable degree of protection against adversarial attacks, despite its misspecification to such opponents.
To further understand why, Table 9 shows the deflection rates of strategic training, measured as the percentage of attacks it is able to prevent out of the successful attacks against an adversarial model:
%$\%\mathrm{deflected}=\frac{\mathrm{strat}(f_{\mathrm{str}})-\mathrm{strat}(f_{\mathrm{adv}})}{\mathrm{clean}(f_{\mathrm{cln}})-\mathrm{strat}(f_{\mathrm{adv}})}$ (11)
@@ 7 Discussion
Acknowledgements
The authors would like to thank Elan Rosenfeld for his helpful discussions. This work is supported by the Israel Science Foundation grant no. 278/22.
@@ Appendix A Broader impact
The field of adversarial learning aims to enable safer deployment of learned models and to make the use of machine learning more secure in real-world applications. The concern regarding attacks on models is concrete, and the possible risks such attacks entail span many domains (Kumar et al., 2020). This underlies the many efforts that are and have been put forth, both in research and in industry, to advance the development of methods for adversarially robust learning.
@@
Our approach to robust learning stems from the observation that in reality, learned models will be susceptible to the influence of the agents who interact with them; and whereas adversarial learning prepares to encounter malicious agents—which are likely the exception—our main thesis is that even non-benign agents are more likely to promote their own interests than to deliberately sabotage learning efforts. As such, we view our work as contributing an economic perspective to how robust learning should be designed, this by acknowledging and modeling the possible utilities of the agents that learning may encounter, and then using this as prior knowledge.
@@
However, as in any domain or task in which agents are modeled, our approach necessitates making assumptions—even if mild—on the nature of their possible incentives. As these are intended to relax the demanding restrictions of adversarial learning, the main limitation of our approach is that the more precisely we model incentives, the more we risk their misspecification. This is an inherent limitation of any method for robust learning: once we commit to an uncertainty set, all guarantees are with respect to this set only. In our work we empirically explore (to some extent) the implications of misspecification (Sec. 6.2). Results are fairly encouraging, and show that quantifying risks can be used as means for the learner to make informed decisions about what learning should seek to be robust against. But safe deployment of strategically trained models in general clearly requires much further investigation of possible failure modes.
Adversarial learning faces similar limitations: it commits to a certain threat model (e.g., $L_{\infty}$ with a certain budget), and provides robustness only against such attacks—but this protection does not generalize to others (e.g., $L_{2}$, or a larger budget) (Hendrycks et al., 2021; Bai et al., 2021). For example, if we train using some attack radius $r$, then we obtain robustness to attacks of at most that strength, and are still susceptible to stronger attacks. One way to think of strategic modeling is in providing more flexibility in designing the uncertainty set by controlling not only the magnitude of attacks (i.e., the attack radius), but also their direction (w.r.t. how the model perceives the ‘directionality’ of classes in input space). Thus, specifying robustness in our framework permits not only attack strength, but also intent.
@@
Nonetheless, in relation to our approach, adversarial learning is indeed ‘safer’, in the sense that for a fixed radius, it protects against all directions, or in our terminology—because it seeks to defend against all possible opponents (i.e, ‘everything’). However, our key point in this paper is precisely that safety that is granted through conservativity comes at a predetermined cost—for example, in reduced clean performance. If we have established knowledge or strong beliefs regarding the opponent’s possible incentives, and are willing to make calculated decisions regarding the gains and risks of making more fine-grained assumptions, then strategic modeling can be highly beneficial as a tool for providing designated robustness against strategic opponents.
@@ B.1 Data and preprocessing
  • CIFAR-10333https://www.cs.toronto.edu/~kriz/cifar.html (Krizhevsky et al., 2014): This dataset includes $60,000$ natural images of size $32\times 32\times 3$. We use the standard split of $50,000$ train and $10,000$ test examples. There are $10$ classes, 4 of which encode vehicles, and 6 of which encode animals. We use this partition as our definition of label semantics.
  • GTSRB [^4] https://benchmark.ini.rub.de (Houben et al., 2013): This dataset includes $\sim 50,000$ images of road signs of variable sizes. We use the standard split of $39,209$ train and $12,630$ test images. For consistency we resize the images to $32\times 32\times 3$. There are $43$ classes, which we partition according to the shape of the sign—either circular or triangular, which roughly corresponds to road sign types: prohibitory and mandatory signs vs. warning signs.
@@ B.2 Architectures
We experiment with three different model architectures: (i) VGG (Simonyan & Zisserman, 2015), (ii) ResNet18 (He et al., 2016), and (iii) ViT (Dosovitskiy et al., 2021). The former two are convolution-based, for which we use standard architectures. The latter is transformer-based, and for this we use an implementation adjusted to smaller resolution images. [^5] https://github.com/omihub777/ViT-CIFAR
@@ B.4 Training and hyperparameters
As we discuss in Sec. 5.1, we found that strategic training benefits (mostly in terms of clean accuracy) from adding subtle noise to the choice of utility at each step. Accordingly, at each batch, for each class $y$, with probability $1-\epsilon$ we use the utility $u_{y\cdot}$ defined in the objective, and with probability $\epsilon$ we draw a target label ${\bar{y}}\in[K]$ uniformly at random. Throughout all experiments, separate from Sec. C.4, we use $\epsilon=0.1$.
@@ B.5 Compute and runtime
Appendix C Additional experiments
C.1 Strategic learning vs. benchmarks against 1-hot opponents
We compare the performance of strategic training to adversarial training and clean training, and respectively measure clean accuracy, strategic accuracy, and adversarial accuracy. In this experiment the strategic opponents are random 1-hot opponents whose utility is known to the learner at test time. Table 2 shows mean accuracies and standard errors for these settings. Results are reported for CIFAR-10 and CIFAR-100 using both VGG and ResNet18 models.
As for ViT, here and in other experiments we found it difficult to obtain comparable performance—even in the clean setting—while remaining within the general regime of the experimental setup of all other experiments (for similar indications, see Mo et al. (2022)).
Results.
As expected, each approach is best on the setting for which it is designed for. In terms of strategic accuracy, strategic training improves significantly over adversarial training, especially in CIFAR-100 ($+30.2$ for VGG, $+25.9$ for ResNet18) but also in CIFAR-10 ($+8.5$ and $+8$). Strategic training also provides a considerable boost in clean accuracy compared to adversarial training. Meanwhile, whereas clean training breaks completely under adversarial evaluation, strategic training still provides some degree of protection against adversarial attacks (between $-9.8$ and $-18.2$).
Table 2: Accuracies under clean, strategic, and adversarial training for random 1-hot opponents.
C.2 Risks and mitigation of imprecise strategic modeling for $k>1$
This experiment complements Sec. 6.2 and measures accuracy for strategic training with increasing risk mitigation $\epsilon$ against $k$-hot opponents with $k>1$ (we show $k=1$ here as well for comparison). As before, results compare performance of strategic training against random opponents (mean and standard deviation; left) as well as against easy (center) and hard (right) opponents, chosen from a large random sample of opponents. We consider both the ‘correct’ setting in which the opponent’s utility is well-specified in the training objective, and in the ‘wrong’ setting in which the objective is misspecified, for which we show performance under the worst-case opponent relative to the trained model. When increasing $\epsilon$, in the well-sepcified setting we expect accuracy to decrease, and in the misspecified setting we expect accuracy to increase.
Results.
Fig. 10 shows that for $k>1$, similar trends are preserved as for $k=1$. One difference is that as $k$ increases, overall performance decreases, as can be expected given that larger $k$ means generally stronger opponents. In the correct setting, note how for the easy opponent the drop in accuracy as a result of increasing $k$ is much smaller than for random or for hard. This suggests that there are inherently ‘easy’ targets that are captured by the easiest (and weakest) opponent having $k=1$. Contrarily, when considering hard opponents, accuracy decreases for $k=1,..,3$ but does not further deteriorate for $k=4$. This suggests that there are inherently ‘hard’ targets, i.e., that the targets encoded in the hardest 3-hot opponent capture most of the difficulty in protecting against stronger opponents. such as those with dominating utilities, $u^{\prime}\geq u$. This implies that defending against these targets alone—rather than against all targets as in adversarial training—suffices to provide reasonable protection in this setting even when utilities are inherently misspecified.
Figure 10: Accuracy for strategic training with increasing risk mitigation ($\epsilon$) for $k\geq 1$.
C.3 Robustness of strategic training against an all-knowing adversary
Our results throughout Sec. 6 show that across multiple settings, strategic learning provides reasonable protection even against adversarial attacks. These results serve to quantify the worst-case risk (in terms of loss in accuracy) a learner can expect by committing to a certain strategic uncertainty set that eventually proves to be erroneous. However, one can also consider an even stronger adversarial model in which the adversary not only attacks the worst-case target, but also knows that the learner is strategic, and hence exploits this to target classes that are left ‘undefended’ by the uncertainty set. We refer to this type of opponent as an all-knowing adversary, and here we evaluate the performance of strategic learning against it.
Setup.
In this experiment we use VGG on CIFAR-10, and consider strategic learning in both the semantic (sem) and anti-semantic (a-sem) settings. In addition to standard measures, here we evaluate performance also w.r.t. the following opponents:
  • avg (a-)sem. 1-hot: an all-knowing adversary that employs a targeted attack on some undefended target, averaged over all such possible targets.
  • min (a-)sem. 1-hot: an all-knowing adversary that employs the worst-case targeted attack of all undefended targets.
  • (a-)sem. multi-target: an all-knowing adversary which attacks all possible undefended targets simultaneously.
  • min 1-hot: the worst-case 1-hot (i.e., targeted) opponent.
Results.
Table 3 shows the performance of strategic training for both the semantic and anti-semantic settings, evaluated against several variations of an all-knowing adversary who knows which targets are defended in learning, and which are left (in principle) unprotected. For comparison, we also show the performance of clean and adversarially trained models against the same attack types. Results show that despite even though strategic training forfeits explicit defense against certain targets, they are nonetheless not fully susceptible to exploitation, even when the adversary knows the learner’s assumed utility structure. This suggests that attack ‘directions’ are likely not orthogonal, and that protecting against some targets indirectly provides protection also against others.
Table 3: Results for strategic learning (semantic and anti-semantic) against an all-knowing adversary.
CIFAR-10 VGG
train test avg a-sem. 1-hot min a-sem. 1-hot min 1-hot a-sem. multi-target adv sem. multi-target clean
sem. 52.3 49.5 43.7 48.5 33.7 50.9 72.2
adv 61.2 58.5 52 56.8 39.9 46.4 67.7
clean 1 0.4 0.1 0.2 0.6 0.1 89.2
train test avg sem. 1-hot min sem. 1-hot min 1-hot sem. multi-target adv a-sem. multi-target clean
a-sem. 43.6 32.2 32.2 23.8 25 69.8 80.3
adv 58.3 50.7 52 46.4 39.9 56.8 67.7
clean 0.4 0.2 0.1 0.1 0.6 0.2 89.2
C.4 Worst-case strategic training against opponents with preferences
Preserving the general setup of Sec. 6.3, here we present results for worst-case strategic learning against opponents with preference-order utilities. This aims to capture settings in which there is a natural ordering over classes that determines their value for the opponent. Examples include:
  • An online market platform for second-hand items (e.g., eBay), where sellers post pictures of their items, and the system classifies them based on quality and condition. Here the potential opponent is the seller, who would like her item to be classified as high quality, and can manipulate the image to achieve this.
  • Similarly, a platform for selling antiques, vintage items, or rarities, in which the system provides buyers with a recommended price range. Here again the seller would like its item to be classified as being in a higher price range.
  • A face recognition system in a security setting where different employees in a firm have different clearance levels or permissions. Here a malicious employee will try to impersonate another only if the other’s security level is higher than his own.
Our main motivation in exploring this setting is to conceptually move further away from the standard adversarial settings, in which the opponent works ‘against’ the learner, and towards settings in which the opponent acts simply to promote its own self-interest. This is in contrast to e.g., semantic and anti-semantic opponents, which still gain from the learner’s losses, though in a structured manner. Preference-order attacks are in principally different; for example, note that in some cases the opponent can correct predictive errors: if an example’s true label is $y$ but is wrongly classified as ${\hat{y}}$, and if $y$ is ranked higher in the opponent’s preference ordering than ${\hat{y}}$, then the attack ${\hat{y}}\mapsto y$ is both helpful for the opponent and for the learner.
Preference order utilities.
For a given ordering $\preceq$ over classes, a preference order opponent (or an opponent with preferences) has utility given by $u_{\preceq}(y,y^{\prime})=\mathds{1}{\{{y\preceq y^{\prime}}\}}$. I.e., if the true class is $y$, then the opponent gains by shifting the prediction to any class $y^{\prime}$ that is ranked higher. Note that for any given $y$, the utility function $u_{\preceq}$ does not prescribe which targets the opponent will necessarily attack—but rather, only which subset of classes are possible targets. Thus, training under $u_{\preceq}$ provides protection against all opponents who have more fine-grained preferences about specific mappings, such as those that could derive from a utility-maximizing opponent having different costs associated with different targets. One interesting aspect of preference utilities is that the number of possible targets for a class $y$ decreases as the importance of $y$ increases in the ranking. In particular, the lowest-ranked class can be mapped to any class (high uncertainty), whereas the highest-ranked class will not be attacked at all.
From a technical perspective, and relative to all others utilities we have so far considered, the key distinction of preference utilities is that they do not allow transitivity: if an opponent gains from $y\mapsto y^{\prime}$, then we can be certain that instances with $y^{\prime}$ will not be attacked to become $y$. This differs from conventional attacks in which transitivity typically holds, often in relation to class similarity. For example, in adversarial attacks, if most dogs are mapped to cats, then we can also expect most cats to be mapped to dogs. This means that the over-conservativity of adversarial training (as well as some forms of strategic training) will manifest in unnecessarily protecting against bi-directional attacks—which will likely be costly in accuracy.
Setup.
In this experiment we focus on CIFAR-10 and use VGG and ResNet18. We report accuracies (mean and std. dev.) averaged over 10 random instances, where in each instance we randomly draw a global $\preceq$ ordering over classes. For implementing preference attacks, we found it beneficial to optimize a variation of Eq. (9) in which the objective also penalizes non-targets:
${\hat{\delta}}_{u}=\operatorname*{argmax}_{\delta}\,\max_{y^{\prime}\in T}\,{\hat{p}}(y^{\prime}|x+\delta)-\max_{y^{\prime}\notin T}\,{\hat{p}}(y^{\prime}|x+\delta)$ (12)
In terms of training, we found that a higher value of $\epsilon=0.5$ was more useful for stabilizing training (see Sec. 5.1). This contributed mostly to improved clean accuracy, while having little impact on strategic accuracy.
Results.
Table 4 presents accuracies under clean training, adversarial training, and strategic training with an uncertainty set that includes all preference-order utilities. As can be seen, strategic training improves significantly vs. adversarial training in terms of both strategic accuracy and clean accuracy, where the former is more significant when using VGG ($+4.1$) and the latter when using ResNet18 ($+2.5$). Interestingly, and despite the fact that preference-based strategic training defends only against non-transitive utilities, it provides reasonable protection against adversarial attacks (which likely are transitive), where for both models performance drops by roughly $-10$ accuracy points. This reduction is not negligible—but is considerably better than under clean training, for which accuracy is virtually zero. One explanation for this is that while each individual preference utility is non-transitive, strategic learning considers the set of all preference utilities simultaneously, which together can provide protection against transitive attack patterns.
Table 4: Accuracies under clean, strategic, and adversarial training for preference-order opponents.
C.5 Worst-case strategic learning on CIFAR-100
Here we extend our experimental analysis from Sec. 6.3 to include the CIFAR-100 dataset. This dataset is similar to CIFAR-10, but includes 100 classes (instead of 10) and has fewer training examples per class (600, rather than 6000). Classes are partitioned into 20 categories (e.g., flowers, fish, trees), where each category includes 5 relevant classes. These categories provided us with a natural definition of semantic groups. Note that not only does CIFAR-100 include more classes (and therefore the size of the utility matrices is much larger), but since there are 20 categories (and therefore 20 semantic groups), in this experiment the anti-semantic uncertainty set is drastically larger than the semantic uncertainty set. Stated differently, whereas for CIFAR-10 both $u^{\mathrm{sem}}$ and $u^{\mathrm{a{\text{-}}sem}}$ roughly half of the entries are non-zero, in CIFAR-100, $u^{\mathrm{sem}}$ has roughly 5% non-zero entries, whereas $u^{\mathrm{a{\text{-}}sem}}$ is roughly 95% positive, and is hence much closer to $u^{\mathrm{adv}}$. Consequently, the improvement of strategic anti-semantic training over adversarial training is expected to be less significant than for CIFAR-10.
Results.
Table 5 presents results for this setting. As can be seen, overall trends are similar to those in Table 1 for CIFAR-10 and GTSRB, with the exception of VGG in the anti-semantic setting for which strategic training does not improve over adversarial training.
Table 5: Results for worst-case strategic learning (semantic and anti-semantic) on CIFAR-100.
C.6 Inferring utilities
In this experiment we explore the extent to which an opponent’s utility matrix can be inferred from their previous attack data. Since attacks are made in response to some model, we consider data from attacks on: (1) a standard model model trained for clean accuracy, and (2) an adversarially-trained model. For each setting, we evaluate the ability to infer attack targets and reconstruct the utility matrix using two ‘levels’ of observed data: (1) models predictions $y’$ ’ alone, and (2) attack vectors $\delta$.
Setup.
Here we focus on CIFAR-10 with ResNet18  and consider random $k$-hot utilities for $k=1,2,3$. Previous attacks are implemented as multi-targeted attacks, corresponding to $k$-hot utilities. We employ two methods for inferring attacks targets, one per type of observed data:
  1. 1.
    Model predictions. Given access only to the initial model’s ‘dirty’ predictions, $f(x+\delta)=y’$ ’, we estimate the target of an attack is precisely the predicted class. This naïve approach will be correct if the attack succeeded, but may fail if not (and the classifier erred regardless).
  2. 2.
    Attack vectors. With access to the attack vectors, $\delta$, we first simulate targeted attacks $\delta_{t}$ on the source image $x$ for each possible target $t$,and then choose the target $\hat{t}$ having the most similar $\delta_{t}$:
    $\hat{t}=\operatorname*{argmin}\nolimits_{t\in[K]}||\delta_{t}-\delta||_{2},\qquad\delta_{t}=\max\nolimits_{\delta^{\prime}}\ell(f(x+\delta^{\prime}),y,t)$ (13)
    where $\ell(\cdot,\cdot,\cdot)$ is the targeted attack loss.
We report two metrics:
  • % Attacks with correct target prediction - The percentage of attacks where the opponent’s target was correctly inferred (excluding cases where ${\hat{y}}=y$, since $u_{yy}=0$ is assumed).
  • % Matrix entries recovered - The percentage of the opponent’s matrix entries correctly recovered. Here we assume that $k$ is known, and for each source class $y$ select the $k$ targets with the highest frequency as predictions.
Results.
Table 6 presents the results for both settings. In the first setting where the initial models is clean, utilities can be inferred with high accuracy using both methods. In the second setting, where the initial model is robust, the model’s predictions alone are insufficient for accurate inference. However, attack vectors enable effective recovery of the opponent’s utility matrix with high accuracy.
Table 6: Results for utility inference on CIFAR-10.
Interestingly, larger $k$ enables better target prediction, but worse reconstruction of matrix entries. The reason for this is twofold. First, in terms of correctly predicted targets, as $k$ increases, it is more likely that the attack succeeds on one of the relevant targets, and so most predictions match these targets of attack. Conversely, for small $k$, there are few viable targets, and predictions tend to be more dispersed across classes. A second-order effect is that for low $k$, some targets are “hard”, and so most attacks fail, the result of which is that predictions become more distributed across classes. However, when $k$ increases, it is more likely that each class has at least one “easy” target, and so attacks mostly focus on this class—and succeed, which makes them easier to identify.
Second, in terms of matrix reconstruction, the above works in the other direction: as $k$ increases, the number of entries with positive utility are likely to include more “hard” targets. Since attacks will likely succeed for the easy targets, the observed data will include only a small number of attempted attacks on the hard targets, which in turn makes their estimation more difficult. This is precisely the challenge of inferring utilities from revealed preferences.
C.7 Utilities in $[0,1]$
Although we define the universal set of utilities as $\Lambda=[0,1]^{K\times K}$, our experiments so far have in effect considered 0-1 utilities of the form $u_{yy^{\prime}}\in\{0,1\}$. One reason is that strategic training with such $u$ gives also protection against any continuous $u^{\prime}$ with $u^{\prime}\leq u$. However, it is also interesting to consider outcomes for defending against utilities with entries in $[0,1]$.
Let $u\in[0,1]^{K\times K}$, and consider an example $(x,y)$. The following procedure implements an attack that maximizes utility for the opponent, which we refer to as a sequential strategic attack:
  1. 1.
    Sort targets $y’\in[K]$ with strictly positive utility values $u_{yy’}>0$ in decreasing order of utility.
  2. 2.
    Attack all such targets sequentially in this order.
    • if an attack succeeds, apply this attack, and break.
If all attack attempts on strictly positive targets fail, we are free to choose how the attack concede. Here we consider a strategy that defaults to applying the (attempted) attack on the most preferred target, but note that other alternatives are similarly plausible.
One implication of the above is that at test-time, only the order of targets with strictly positive utility matters—not their actual value. Furthermore, if we replace a continuous $u$ with a 0-1 utility matrix $\bar{u}$ such that $\bar{u}_{yy^{\prime}}=1$ iff $u_{yy^{\prime}}>0$, then a sequential attack based on $u$ is equivalent to a multi-targeted attack based on $u^{\prime}$. In this sense, for a fixed model, performance against an opponent with $u$ in $[0,1]$ and an opponent with a matching 0-1 utility $\bar{u}$ is the same.
Nonetheless, there can be a difference in training with $u$ vs. $\bar{u}$, in the sense that strategic training can output different models in each case. Intuitively, if attacks against lower-ordered targets never materialize, then learning might benefit (e.g., in terms of clean accuracy) by reducing the need to defend against such targets.
Setup.
We consider CIFAR-10 with ResNet18, and compare the following training methods:
  • Sequential strategic training – The simulated attacker mimics the described $[0,1]$ attack strategy.
  • Sequential strategic training with adversarial fallback – similar to the above, but in case of failure on all strictly positive targets, applies an adversarial attack. This presents a slightly more conservative approach using the available degree of freedom.
  • Utility-weighted strategic training – a naïve approach which weighs the logits in the attack objectives according to the utility, namely:
    $\delta=\operatorname*{argmax}\nolimits_{y^{\prime}}p(y^{\prime}\mid x+\delta)\cdot u_{yy^{\prime}}$
    Although this appears to be a natural generalization of Eq. (9) from 0-1 utilities to $[0,1]$ utilities, this attack implementation does not maximize utility, since it considers the model’s internal scores—not actual prediction outcomes.
  • Multi-targeted training – uses the strategic attack in Eq. (9) on the matching 0-1 utility matrix $\bar{u}$.
  • Adversarial training as a baseline.
We evaluate these models against $[0,1]$ attackers using two types of 2-hot utilities: (1) Aligned - for each class, the higher-utility target is “easy” to attack, and harder to defend against (e.g., cat $\rightarrow$ dog) compared to the lower-utility target (e.g., cat $\rightarrow$ truck). We refer to this utility as aligned because its ordering corresponds to attack feasibility, matching the targets a multi-targeted attacker is likely to attack. (2) Misaligned - using the same set of targets as in (1), but with the ordering reversed.
Results.
Table 7 presents results for both aligned and misaligned utilities. In the aligned case, differences in performance between the various models against the $[0,1]$ attack are smaller. This supports our conjecture, that when the attacker’s ordering is aligned with attack feasibility, both multi-targeted and sequential training protect against the same targets. Conversely, in the misaligned case, sequential training demonstrates a significant advantage over multi-targeted training, in terms of both $[0,1]$ strategic accuracy and clean accuracy.
Table 7: Results for [0,1] utilities on CIFAR-10.